OpenStudy (anonymous):

prime factorization of 98

OpenStudy (asevilla5):

Break 98 into factors and then, each factor by factor to get.. 98 = 2 * 49 = 2 * 7 * 7 = 2 * 7² I hope this helps!

OpenStudy (mokeira):

basically just use prime numbers to factor out i have used 2 and 7 ...and they are prime numbers :) do you get my point

jimthompson5910 (jim_thompson5910):

I like Mokeira's approach and it's often handy to circle the primes so you can write out the prime factorization